OpenStudy (anonymous):

Use the substitution method to find the solution(s) to the system of equations. Show all work for full credit. y = –x2 y = x

OpenStudy (anonymous):

sub y=x into y=-x^2 x=-x^2 x^2 + x =0 x(x+1) =0 x=0 or -1

OpenStudy (anonymous):

what school do u go to?

OpenStudy (anonymous):

flvs

OpenStudy (anonymous):

same lol no wonder that question is so familiar.

OpenStudy (anonymous):

yeah im taking my test now and its killing me

OpenStudy (anonymous):

im tired with tests, but biology just blows my mind!